# Non-standard Holomorphic Structures on Line Bundles over the Quantum Projective Line

### v1

@@ 1 Introduction
Over the past three decades, noncommutative differential geometry has witnessed substantial progress [connes1994noncommutative]. By contrast, the corresponding theory of noncommutative complex geometry is still at a relatively early stage of development. An important step in this direction was made by Khalkhali, Landi, and Van Suijlekom in [khalkhali2011holomorphic], where the authors introduced, for a deformation parameter $0<q<1$, the quantum projective line ${\mathbb{C}}P^{1}_{q}$. This space provides a rich and instructive example of a noncommutative complex manifold. They further demonstrated that many of the classical features of the complex projective line ${\mathbb{C}}P^{1}$ continue to hold in the quantum setting. In particular, for each $n\in\mathbb{Z}$ they constructed holomorphic line bundles $\mathcal{L}_{n}$ on ${\mathbb{C}}P^{1}_{q}$, which may be regarded as noncommutative analogs of the classical line bundles $\mathcal{O}(n)$ on ${\mathbb{C}}P^{1}$.
@@ Acknowledgments
The authors would like to thank Jonathan Block for his encouragement and valuable suggestions. They are also grateful to David Gao and Réamonn Ó Buachalla for kindly answering questions related to this work, and to Ryszard Szwarc for directing them to the reference [koekoek2010hypergeometric].
@@ 2.3 $C^{*}$-completions, $L^{2}$-completions, and the spectral triple
[[das2020dolbeault] Theorem 6.21] $(\mathcal{A}({\mathbb{C}}P^{1}_{q}),L^{2}(\Omega^{0,\bullet}({\mathbb{C}}P^{1}_{q})),D_{\bar{\partial}})$ forms a spectral triple in the sense of [connes1994noncommutative], where $D_{\bar{\partial}}:=\bar{\partial}+\bar{\partial}^{\dagger}$.
@@ 3.1 Generalities on gauge equivalences of holomorphic structures
$\bar{\partial}(B_{0}^{n})=\sum_{k=0}^{n-1}q^{2k}B_{0}^{n-1}\bar{\partial}(B_{0})=q^{n-1}[n]_{q^{2}}B_{0}^{n-1}\bar{\partial}(B_{0}),$ (3.1)
where $n]_{q^{2}}=\frac{q^{2n}-q^{-2n}}{q^{2}-q^{-2}}$ is the $q^{2}$-integer as in [khalkhali2011holomorphic, (3.1)]. We define $n]_{q^{2}}!:=\prod_{k=1}^{n}[k]_{q^{2}}$ and $0]_{q^{2}}!:=1$. Then we can check that
$g:=\sum_{n=0}^{\infty}\frac{B_{0}^{n}}{q^{\frac{n(n-1)}{2}}[n]_{q^{2}}!}$ (3.2)
satisfies $\bar{\partial}(g)=g\bar{\partial}(B_{0})$. It is clear that $g\in\mathcal{C}({\mathbb{C}}P^{1}_{q})$ is invertible and $g\in{\mathrm{Dom}}(\bar{\partial})$ but $g\notin\mathcal{A}({\mathbb{C}}P^{1}_{q})$.
Inspired by Example Example we have the following modified definition.
@@
We call two $\bar{\partial}$-connections $\overline{\nabla}_{\theta_{1}},\overline{\nabla}_{\theta_{2}}$ on $\mathcal{L}_{n}$ gauge equivalent if there exists an invertible element $g\in\mathcal{C}({\mathbb{C}}P^{1}_{q})^{\times}\cap{\mathrm{Dom}}(\bar{\partial})$ such that
@@
In particular, a $\bar{\partial}$-connection $\overline{\nabla}_{\theta}$ on $\mathcal{L}_{n}$ is gauge equivalent to the standard $\bar{\partial}$-connection if there exists an invertible element $g\in\mathcal{C}({\mathbb{C}}P^{1}_{q})^{\times}\cap{\mathrm{Dom}}(\bar{\partial})$ such that
@@
By Proposition 2.10 we know that $H^{0,1}_{\overline{\nabla}}(\mathcal{L}_{0})=0.$ As a result for any $\theta\in\Omega^{0,1}({\mathbb{C}}P^{1}_{q})$ there exists an $f\in\mathcal{A}({\mathbb{C}}P^{1}_{q})$ such that $\theta=\bar{\partial}(f)$. Therefore a $\bar{\partial}$-connection $\overline{\nabla}_{\theta}=\overline{\nabla}_{\bar{\partial}(f)}$ on $\mathcal{L}_{n}$ is gauge equivalent to the standard $\bar{\partial}$-connection if there exists an invertible element $g\in\mathcal{C}({\mathbb{C}}P^{1}_{q})^{\times}\cap{\mathrm{Dom}}(\bar{\partial})$ such that
@@
If there exists a non-zero non-invertible $h\in\mathcal{C}({\mathbb{C}}P^{1}_{q})\cap{\mathrm{Dom}}(\bar{\partial})$ such that $\bar{\partial}(h)=h\bar{\partial}(f)$, then there cannot exist an invertible $g$ such that $\bar{\partial}(g)=g\bar{\partial}(f)$.
@@
For the sake of contradiction, let such a $g$ exist. Since $g$ is invertible, we can write $h=ag$ for $a=hg^{-1}$. We have
@@ 3.2 The $C^{*}$-subalgebra $C^{*}(1,B_{0})$
where $C(sp(B_{0}))$ is the $C^{*}$-algebra of continuous functions on $sp(B_{0})$ the spectrum of $B_{0}$. Recall Proposition 2.16 tells us
@@ 3.3 Existence of non-standard holomorphic structures
Given $f\in\mathcal{A}({\mathbb{C}}P^{1}_{q})\cap C^{*}(1,B_{0}),$ there exists an invertible $g\in\mathcal{C}({\mathbb{C}}P^{1}_{q})^{\times}\cap{\mathrm{Dom}}(\bar{\partial})$ such that $\bar{\partial}g=g\bar{\partial}f$ if and only if $f$ has no defective spot.
@@
By Corollary 3.22, if $f$ has no defective spot. an invertible solution $g$ to $\bar{\partial}g=g\bar{\partial}f$ exists.
On the other hand, if $f$ has a defective spot, then by Corollary 3.22, $\bar{\partial}g=g\bar{\partial}f$ has a not-invertible, nonzero solution. By Lemma 3.5, $\bar{\partial}g=g\bar{\partial}f$ cannot have any invertible solution. ∎
Example.
We notice that the element $B_{0}$ has no defect spot. Actually in Example Example we found explicitly an invertible element $g$ such that $\bar{\partial}g=g\bar{\partial}B_{0}$.
@@ 3.4 Gauge equivalence between $\bar{\partial}$-connections
We proved that (3.27) holds for any monomials hence for any polynomials. The general case now follows from the fact that $\bar{\partial}$ is a closed operator and multiplication and $m_{q^{2}}$ are bounded operators. ∎
@@
The second assertion follows from the first one and Definition 3.2.
We again use the functional calculus isomorphism $\Psi$ to identify $C^{*}(1,B_{0})$ and $C(sp(B_{0}))$. Equation (3.31) then becomes
@@ 5 Future Work
We also notice that higher dimensional quantum projective spaces ${\mathbb{C}}P^{l}_{q}$ and line bundles over them were introduced and studied in [khalkhali2011noncommutative]. It is interesting to study non-standard holomorphic structures on line bundles over ${\mathbb{C}}P^{l}_{q}$. The analysis will be more involved in higher dimensional case as the flatness condition $\overline{\nabla}_{\theta}\circ\overline{\nabla}_{\theta}=0$ does not hold automatically on ${\mathbb{C}}P^{l}_{q}$ for $l\geq 2$.

### v2

@@ 1 Introduction
Over the past three decades, noncommutative differential geometry has witnessed substantial progress [connes1994noncommutative]. By contrast, the corresponding theory of noncommutative complex geometry is still at a relatively early stage of development. An important step in this direction was made by D’Andrea, Dabrowski, and Landi in [d2008noncommutative], where the authors introduced, for a deformation parameter $0<q<1$, the quantum projective plane ${\mathbb{C}}P^{2}_{q}$. These spaces provide rich and instructive examples of noncommutative complex manifolds. For higher dimensional cases and further discussion, see [d2010dirac], [d2010anti], and [d2013geometry]. Khalkhali, Landi, and Van Suijlekom in [khalkhali2011holomorphic] studied holomorphic structures on the quantum projective line ${\mathbb{C}}P^{1}_{q}$ in details. They demonstrated that many of the classical features of the complex projective line ${\mathbb{C}}P^{1}$ continue to hold in the quantum setting. In particular, for each $n\in\mathbb{Z}$ they constructed holomorphic line bundles $\mathcal{L}_{n}$ on ${\mathbb{C}}P^{1}_{q}$, which may be regarded as noncommutative analogs of the classical line bundles $\mathcal{O}(n)$ on ${\mathbb{C}}P^{1}$.
@@ Acknowledgments
The authors would like to thank Jonathan Block for his encouragement and valuable suggestions. They are also grateful to David Gao and Réamonn Ó Buachalla for kindly answering questions related to this work, to Ryszard Szwarc for directing them to the reference [koekoek2010hypergeometric]. They want to thank Francesco D’Andrea for helpful comments on this work and Satyajit Guin for pointing out a mistake in the first version of this paper.
@@ 2.3 $C^{*}$-completions, $L^{2}$-completions, and the spectral triple
[[d2010dirac] Theorem 6.2, see also [das2020dolbeault] Theorem 6.21] $(\mathcal{A}({\mathbb{C}}P^{1}_{q}),L^{2}(\Omega^{0,\bullet}({\mathbb{C}}P^{1}_{q})),D_{\bar{\partial}})$ forms a spectral triple in the sense of [connes1995noncommutative], where $D_{\bar{\partial}}:=\bar{\partial}+\bar{\partial}^{\dagger}$.
@@ 3.1 Generalities on gauge equivalences of holomorphic structures
$\bar{\partial}(B_{0}^{n})=\sum_{k=0}^{n-1}q^{2k}B_{0}^{n-1}\bar{\partial}(B_{0})=q^{n-1}[n]_{q}B_{0}^{n-1}\bar{\partial}(B_{0}),$ (3.1)
where $n]_{q}=\frac{q^{n}-q^{-n}}{q-q^{-1}}$ is the $q$-integer as in [khalkhali2011holomorphic, (3.1)]. We define $n]_{q}!:=\prod_{k=1}^{n}[k]_{q}$ and $0]_{q}!:=1$. Then we can check that
$g:=\sum_{n=0}^{\infty}\frac{B_{0}^{n}}{q^{\frac{n(n-1)}{2}}[n]_{q}!}$ (3.2)
satisfies $\bar{\partial}(g)=g\bar{\partial}(B_{0})$. Notice that Propositioin 2.16 gives us $\lVert B_{0}\rVert=1$ hence the series in (3.2) convergies. It is clear that
$g^{-1}=\sum_{n=0}^{\infty}\frac{(-1)^{n}B_{0}^{n}}{q^{\frac{n(n-1)}{2}}[n]_{q}!}$
hence $g\in\mathcal{C}({\mathbb{C}}P^{1}_{q})$ is invertible and $g\in{\mathrm{Dom}}(\bar{\partial})$ and $g^{-1}\in{\mathrm{Dom}}(\bar{\partial})$ but $g\notin\mathcal{A}({\mathbb{C}}P^{1}_{q})$.
Inspired by Example 1 we have the following modified definition.
@@
We call two $\bar{\partial}$-connections $\overline{\nabla}_{\theta_{1}},\overline{\nabla}_{\theta_{2}}$ on $\mathcal{L}_{n}$ gauge equivalent if there exists an invertible element $g\in\mathcal{C}({\mathbb{C}}P^{1}_{q})^{\times}\cap{\mathrm{Dom}}(\bar{\partial})$ such that $g^{-1}\in{\mathrm{Dom}}(\bar{\partial})$ and
@@
In particular, a $\bar{\partial}$-connection $\overline{\nabla}_{\theta}$ on $\mathcal{L}_{n}$ is gauge equivalent to the standard $\bar{\partial}$-connection if there exists an invertible element $g\in\mathcal{C}({\mathbb{C}}P^{1}_{q})^{\times}\cap{\mathrm{Dom}}(\bar{\partial})$ such that $g^{-1}\in{\mathrm{Dom}}(\bar{\partial})$ and
@@
By Proposition 2.10 we know that $H^{0,1}_{\overline{\nabla}}(\mathcal{L}_{0})=0.$ As a result for any $\theta\in\Omega^{0,1}({\mathbb{C}}P^{1}_{q})$ there exists an $f\in\mathcal{A}({\mathbb{C}}P^{1}_{q})$ such that $\theta=\bar{\partial}(f)$. Therefore a $\bar{\partial}$-connection $\overline{\nabla}_{\theta}=\overline{\nabla}_{\bar{\partial}(f)}$ on $\mathcal{L}_{n}$ is gauge equivalent to the standard $\bar{\partial}$-connection if there exists an invertible element $g\in\mathcal{C}({\mathbb{C}}P^{1}_{q})^{\times}\cap{\mathrm{Dom}}(\bar{\partial})$ such that $g^{-1}\in{\mathrm{Dom}}(\bar{\partial})$ and
@@
Although it seems plausible that $g\in\mathcal{C}({\mathbb{C}}P^{1}_{q})^{\times}\cap{\mathrm{Dom}}(\bar{\partial})$ implies $g^{-1}\in{\mathrm{Dom}}(\bar{\partial})$, we do not know whether it is true in general. So we have to add $g^{-1}\in{\mathrm{Dom}}(\bar{\partial})$ to the definition. The authors want to thank Satyajit Guin for pointing out this.
Also see Corollary 3.21 below for a proof in a special case.
Remark 3.4.
@@
If there exists a non-zero non-invertible $h\in\mathcal{C}({\mathbb{C}}P^{1}_{q})\cap{\mathrm{Dom}}(\bar{\partial})$ such that $\bar{\partial}(h)=h\bar{\partial}(f)$, then there cannot exist an invertible $g$ such that $g\in\mathcal{C}({\mathbb{C}}P^{1}_{q})^{\times}\cap{\mathrm{Dom}}(\bar{\partial})$, and $g^{-1}\in{\mathrm{Dom}}(\bar{\partial})$, and $\bar{\partial}(g)=g\bar{\partial}(f)$.
@@
For the sake of contradiction, let such a $g$ exist. Since $g$ is invertible, we can write $h=ag$ for $a=hg^{-1}$. Since both $h$ and $g^{-1}$ are in ${\mathrm{Dom}}(\bar{\partial})$, so is $a$. We then have
@@ 3.2 The $C^{*}$-subalgebra $C^{*}(1,B_{0})$
where $C(sp(B_{0}))$ is the $C^{*}$-algebra of continuous functions on $sp(B_{0})$ the spectrum of $B_{0}$. In particular $\Psi$ maps $B_{0}$ to the indentity function $f(x)=x$ on $p(B_{0})$. Recall Proposition 2.16 tells us
@@
If $f\in C(sp(B_{0}))$ is in the domain of $\bar{\delta}$ and $f$ is invertible, then $f^{-1}$ is also in the domain of $\bar{\delta}$.
Proof.
From the definition, we know that $f\in C(sp(B_{0}))$ is invertible if and only if $f(q^{2n})\neq 0$ for each $n\geq 0$ and $f(0)=\lim_{n\to\infty}f(q^{2n})\neq 0$. If we further know that $f$ is in the domain of $\bar{\delta}$, then by Proposition 3.20 we have
$\lim_{k\to\infty}\frac{f(q^{2k})-f(q^{2k+2})}{q^{2k}-q^{2k+2}}=\text{ exists.}$
Therefore
$\begin{split}&\lim_{k\to\infty}\frac{f^{-1}(q^{2k})-f^{-1}(q^{2k+2})}{q^{2k}-q^{2k+2}}\\ =&-\lim_{k\to\infty}\frac{1}{f(q^{2k})f(q^{2k+2})}\lim_{k\to\infty}\frac{f(q^{2k})-f(q^{2k+2})}{q^{2k}-q^{2k+2}}\text{ also exists.}\end{split}$
Again by Proposition 3.20 we know that $f^{-1}$ is in the domain of $\bar{\delta}$. ∎
Corollary 3.22.
@@ 3.3 Existence of non-standard holomorphic structures
Given $f\in\mathcal{A}({\mathbb{C}}P^{1}_{q})\cap C^{*}(1,B_{0}),$ there exists an invertible $g\in\mathcal{C}({\mathbb{C}}P^{1}_{q})^{\times}\cap{\mathrm{Dom}}(\bar{\partial})$ such that $g^{-1}\in{\mathrm{Dom}}(\bar{\partial})$ and $\bar{\partial}g=g\bar{\partial}f$ if and only if $f$ has no defective spot.
@@
By Corollary 3.22 and Corollary 3.24, if $f$ has no defective spot, then an invertible solution $g\in\mathcal{C}({\mathbb{C}}P^{1}_{q})^{\times}\cap C^{*}(1,B_{0})\cap{\mathrm{Dom}}(\bar{\partial})$ to $\bar{\partial}g=g\bar{\partial}f$ exists. By Corollary 3.21 and Corollary 3.22, we know that $g^{-1}\in{\mathrm{Dom}}(\bar{\partial})$.
On the other hand, if $f$ has a defective spot, then by Corollary 3.24, $\bar{\partial}g=g\bar{\partial}f$ has a non-invertible, nonzero solution. By Lemma 3.6, $\bar{\partial}g=g\bar{\partial}f$ cannot have any invertible solution with $g$ and $g^{-1}$ in ${\mathrm{Dom}}(\bar{\partial})$. ∎
Example 2.
We notice that the element $B_{0}$ has no defect spot. Actually in Example 1 we found explicitly an invertible element $g$ such that $g^{-1}\in{\mathrm{Dom}}(\bar{\partial})$ and $\bar{\partial}g=g\bar{\partial}B_{0}$.
@@ 3.4 Gauge equivalence between $\bar{\partial}$-connections
We proved that (3.27) holds for any monomials hence for any polynomials. The general case now follows from the fact that $\bar{\partial}$ is a closed operator, together with the fact that multiplication and $m_{q^{2}}$ are bounded operators. ∎
@@
The second assertion follows from the first one together with Definition 3.2 and Corollary 3.21.
To prove the first assertion, we again use the functional calculus isomorphism $\Psi$ to identify $C^{*}(1,B_{0})$ and $C(sp(B_{0}))$. Equation (3.31) then becomes
@@ 5 Future Work
We also notice that higher dimensional quantum projective spaces ${\mathbb{C}}P^{l}_{q}$ and line bundles over them were studied in [d2008noncommutative], [d2010anti], [d2010dirac], [d2013geometry], and [khalkhali2011noncommutative]. It is interesting to study non-standard holomorphic structures on line bundles over ${\mathbb{C}}P^{l}_{q}$. The analysis will be more involved in higher dimensional case as the flatness condition $\overline{\nabla}_{\theta}\circ\overline{\nabla}_{\theta}=0$ does not hold automatically on ${\mathbb{C}}P^{l}_{q}$ for $l\geq 2$.